# Foregrounding Artist Opinions: A Survey Study on Transparency, Ownership, and Fairness in AI Generative Art

### 30 Jan 2024 v2

@@ Abstract.
Generative Artificial Intelligence (AI) tools are used to create art-like outputs and aid in the creative process. While these tools have potential benefits for artists, they also have the potential to harm the art workforce and infringe upon artistic and intellectual property rights. Without explicit consent from artists, Generative AI creators scrape artists’ digital work to train Generative AI models and produce art-like model outputs at scale. These outputs are now being used to compete with human artists in the marketplace as well as being used by some artists in their generative processes to create art. We surveyed 459 artists to investigate the tension between artists’ opinions on Generative AI art’s potential utility and harm. This study surveys artists’ opinions on the utility and threat of Generative AI art models, fair practices in the disclosure of artistic works in AI art training models, ownership and rights of AI art derivatives, and fair compensation. We find that artists, by and large, think that model creators should be required to disclose in detail what art and images they use to train their AI models. We also find that artists’ opinions vary by professional status and practice, demographics, whether they have purchased art, and familiarity with and use of Generative AI. We hope the results of this work will further more meaningful collaboration and alignment between the art community and Generative AI researchers and developers.
AI Art, Generative AI, AI Ethics, Survey Study, Consent, Privacy
@@ 1. Introduction
Art has a rich history in co-evolution with the development of new technologies (Etro and Galenson, 2023). The advent of Artificial Intelligence (AI) tools used to create or assist in the creative process carries great potential utility for artists but also harbors the significant potential for harm to the art workforce and artistic and intellectual property. To examine this tension between utility and threat, we surveyed 459 artists about their views on AI-generated art.
@@
We summarize our research results as follows: (Q1) 61.87% of survey participants agree that AI models are a threat to art workers. (Q2) 44.88% of survey participants agreed that AI art models are a positive development in the field of art (22.03% of survey participants agreed that both AI was a positive development and a threat). (Q3) Artists overwhelmingly agreed (80.17%) that model creators should be required to disclose in detail what art and images they use to train their AI models. (Q4) Artists were relatively split between thinking that derivatives from AI art models should be owned by the original artists whose work is represented in the AI art (41.39%) and the user who used the model to create the AI art (39.22%). In our study, 26.80% of the participants agreed that the work and its derivatives be considered the property of the AI model creators. We also looked into what types of compensation artists consider to be fair in exchange for the use of their artwork to train an AI art model. Half of the respondents (50.97%) stated that they did not need compensation, but what mattered to them was who was profiting from the use of their artwork. The most common fair compensation option selected by the survey participants was “I don’t need to profit, but I don’t want for-profit companies to profit from my art” (22.80%). This suggests that artists are not entirely against their art being used in AI art model training data but oppose for-profit companies benefiting financially. We describe the survey results in detail in Section 3 and determine which characteristic variables of our survey participants are most associated with these results. We will present the survey methodology in Section A.1.
@@ 2. Background
There is ongoing disagreement regarding whether the output of Generative AI can be considered art (AI, 2023a; Commission, 2023; Jiang et al., 2023); we will not examine the topic in depth here. For the sake of convenience, we will refer to the comparable output of AI as “art” throughout this paper. However, this is not intended to refute (or endorse) the claim that AI cannot create art.
According to work by Jiang et al. (Jiang et al., 2023), “following philosophers of art and aesthetics from varied disciplines… we define art as a uniquely human endeavor connected specifically to human culture and experience.” They distinguish between inspiration/training and technique/creation. They also point out that a person being able to mimic another person’s artistic style so perfectly as to be indistinguishable is rare, as good forgeries are typically noteworthy. They claim that these aspects, among others, differentiate between what humans and AI can create.
Another dimension of debate concerning whether Generative AI creates art is whether the output of the model is copyrightable under US law. The Courts in the United States recently rejected for copyright a work the plaintiff described as “autonomously created by a computer algorithm running on a machine,” saying there is a “consistent understanding that human creativity is the sine qua non at the core of copyrightability, even as that human creativity is channeled through new tools or into new media” (US District Court for the District of Columbia, 2023).
Among the arguments intended to distinguish human art from AI-generated art, we find three that are particularly compelling. First, human art is the product of activity without “organic pressure” – allowing us to distinguish between the music of people, songbirds, and AI models (Jiang et al., 2023). Second, that the world and the artist are always tightly coupled (Jiang et al., 2023), in part through corporeality (Pfeifer and Bongard, 2006). Third, that art plays a special role in the lived experience of humans (Huyghe, 1970).
Regardless of whether AI can make art, the output of Generative AI can meet the minimally viable criteria for a commercial product, like the cover of a book (Crabapple, 2022) (see Sec. 2.3), which means it is making its way into some of the many places art currently lives in our world, impacting artists and others along the way.
@@
It is important to note that, although how technology can be incorporated into art may be up for debate, that technology is part of the artistic process is not new or ambiguous. Artists have always used tools (Huyghe, 1970), and many art forms involve machines (such as woodwork, metalwork, and digital art) and automation (such as digital printers and cameras). Some artists also choose to use technology in ways other than its intended purpose or develop new art-making techniques by breaking technology. Such artists may be both suspicious and intrigued by the technology they use, questioning the relationship between humans and machines and expanding art practices and processes. When presented with new technology, an artist may ask questions like “How can I keep my work with the machine intentional? How might this technology distract me? Is the machine using me, or am I using it? Is this image-making or art-making? How does this automation serve my artistic intentions and practice?” (Karson, 2023).
@@
In the case of machine learning, a crucial question for many artists is “What data was used to train this model?” (See Sec. 3.) Such inquiries are part of the artistic tradition of challenging the status quo and are in stark contrast to how the creators of Generative AI models have presented their tools for use to the general public (AI, 2023a).
2.1. Generative AI
The field of Generative AI has created much excitement around its relevant applications, both anticipated and emergent (Harby, 2023; Bubeck et al., 2023; Abdin et al., 2023). Its valuation of tens of billions of dollars (Jiang et al., 2023; for Human-Centered Artificial Intelligence, 2023; Crabapple, 2022) has, of course, inspired a gold rush of research (academic, industrial, and otherwise) and media coverage (Shan et al., 2023; Gozalo-Brizuela and Garrido-Merchán, 2023). Many thousands of models are available and accessible (Face, 2023), along with an ecosystem of tools supporting their use. (Gozalo-Brizuela and Garrido-Merchán, 2023) The space of related work being explored, therefore, extends from well-established benchmarking (Abdin et al., 2023; Lynch, 2023; for Human-Centered Artificial Intelligence, 2023), all the way to cynical (Nicoletti and Bass, 2023), often predatory (Hall, 2023), money-making schemes and fraud (Horton, 2023). Since 2010, the number of yearly AI-related publications has only grown, reaching nearly 500,000 in 2021 alone (for Human-Centered Artificial Intelligence, 2023).
The core functionality of a Generative AI model is as follows: based on input from a person – through a user interface, as a conversational text prompt, or as a piece of media to use as a starting point, such as a video, text, code snippet, or an image – Generative AI models can produce novel output in an equal variety of media (see Fig. 2(Gozalo-Brizuela and Garrido-Merchán, 2023).
Currently, the major players in textual Generative AI are large language models (LLMs), such as OpenAI’s ChatGPT and GPT-4 (Zhao et al., 2023; AI, 2023b). These models work by learning how to provide a plausible next word in response to a sequence of text, which results in seemingly fluent linguistic output (Zimmerman et al., 2023). In the visual sphere, models like StableDiffusion and Midjourney (Holz and team, 2024) can modify an image by iteratively tweaking random noise until they create an image that matches the user’s input. The model decides which part of the image needs to be modified at any given time based on the diffusion process (Andrew, 2023).
@@ 2.1. Generative AI
For the purposes of our paper, we define Generative AI as being a subset of tools or processes that combine massive probability distributions with sense-making algorithms in order to create output that, taken as a whole, was not included in the training data (Jiang et al., 2023). The probability distribution should be based on the training data. The sense-making algorithm could be based on several strategies, such as a physical process like diffusion or an architectural approach like transformers/neural nets, although in the end, it should produce something akin to a latent/embedding space (Shan et al., 2023); a place to hold meaning (Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023; Grand et al., 2022). Research is ongoing into the various distributions and processes that can be implemented in Generative AI models to achieve specific behaviors and results (Xu et al., 2023). However, irrespective of details, all of these models use advanced architectures in conjunction with large amounts of data gathered through brute-force methods across vast swathes of the publicly accessible internet  (Xu et al., 2023; Zhao et al., 2023; Salkowitz, 2022).
Figure 2. Generative AI models can produce novel output in various media such as video, text, code snippet, or image. This output is based on the input from an individual through a user interface, conversational text prompt, or a piece of media, as mentioned in a survey conducted by Gozalo Brizuela (Gozalo-Brizuela and Garrido-Merchán, 2023). The technology is advancing rapidly, and in many contexts, the output generated by these models is indistinguishable from that of a human.
A schematic of typical inputs to Generative AI models such as text and images and the resulting outputs (e.g., text, images, video, code, music)
Figure 2. Generative AI models can produce novel output in various media such as video, text, code snippet, or image. This output is based on the input from an individual through a user interface, conversational text prompt, or a piece of media, as mentioned in a survey conducted by Gozalo Brizuela (Gozalo-Brizuela and Garrido-Merchán, 2023). The technology is advancing rapidly, and in many contexts, the output generated by these models is indistinguishable from that of a human.
@@
Here, we are interested specifically in how artists perceive Generative AI, what its impact on art-related careers might be, and what moral, ethical, artistic, and financial implications are coming hand-in-hand with the broader flourishing of Generative AI (Shan et al., 2023; Baio, 2022; Crabapple, 2022; Murphy, 2022; for Human-Centered Artificial Intelligence, 2023; Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023). We, therefore, touch on four relevant topics: first, related academic work; second, the economic context for this discussion; third, what sort of stages are within the pipeline of Generative AI models and how those stages can each cause harm (a specific and noteworthy type of impact); fourth, an example of how a previous technological change caused upheaval within a creative field.
@@
Although in our paper we focus more on the pressing potential for harm, it is worth noting that there is also growing excitement around Generative AI’s capabilities and how it could be incorporated into artistic processes (Karson, 2023). For example, artist Cosmo Wenman worked with blind researchers Brandon Biggs, Lindsay Yazzolino, and Joshua Miele to create images using the Midjourney tool (Holz and team, 2024). Their team noted that programs like Midjourney can find “weird and thematic intersections” in art that go unnoticed by human eyes. They also discussed the role of chaos in creating “visual intrigue” and how technology can be used to “enhance agency.” Wenman excitedly called the new technology “essentially moveable type for visual rhetoric” (Schrader, 2023).
@@ 2.2. Related work
Jiang et al.(Jiang et al., 2023) analyze Generative AI art models and their potential impact on the world through the lens of various academic sources (both philosophical and scientific). They describe many opportunities for harm as a result of such products: violation of the consent and privacy of artists (Crabapple, 2022; Murphy, 2022; Baio, 2022; Ho, 2024; the Midjourney team, 2023), a lack of dataset transparency (shielding violations and biases) (Ho, 2024; the Midjourney team, 2023), a failure to uphold the contextual integrity of content (Nissenbaum, 2004), harm to the livelihood of the artist, biases in the training data, automating processes in a way that is detrimental to the creation of art, reduced accessibility, erasure of human works by obscurity, and increasing reluctance by artists to share their works openly (Crabapple, 2022).
The harms Jiang et al. (Jiang et al., 2023) foresee for artists fall into two categories, financial and emotional, which manifest primarily as the replacement of a human artist on a project, the impersonation of an artist without consent (Baio, 2022; Murphy, 2022; Crabapple, 2022; Gozalo-Brizuela and Garrido-Merchán, 2023), or plagiarism. Impersonation, particularly in this context, could be weaponized for nefarious purposes (Horton, 2023; Ho, 2024; the Midjourney team, 2023). Moreover, these two kinds of harm, enacted against individuals at scale, create societal-level harm. Specific societal harms Jiang et al. discuss in their work include data laundering – through academic or open-source projects where public funds are funneled into research with ultimately private profits – the proliferation of stereotypes and “hegemonic views,” and erasure by obscurity (i.e., that AI-generated outputs can be created much more quickly than human ones). They also point out that US copyright and intellectual property jurisprudence needs to be better equipped to deal with these new tools and that open sourcing, often called on as a solution to software-related problems, is incomplete in this case (Jiang et al., 2023).
@@ 2.3. Economic context: livelihood of the artist
AI (a term which includes, but is not limited to, Generative AI) is increasingly working its way into the world economy. Across nearly every economic sector for which data is gathered, AI-related job postings have been increasing (for Human-Centered Artificial Intelligence, 2023). Globally, employers in a broad range of fields seek AI-skilled workers, and every year, many companies decide to adopt AI tools. The top benefit these companies claim to see in response to that decision is lowered costs (for Human-Centered Artificial Intelligence, 2023), including through layoffs and replacing employees with AI (Korn, 2024). On the other end of the employment ladder, workers in an increasing variety of fields list AI as a relevant skill (for Human-Centered Artificial Intelligence, 2023).
Some of these impacts are already coming to bear on the economic reality of artists. Entry-level positions in film, TV, and gaming seem particularly at risk, and the conceptual period of a project seems to be a popular phase in which to replace artists with AI (Carter, 2023). Traditional niches like illustrating book covers can now be filled by the output of Generative AI (Crabapple, 2022; Chalk, 2024), and major employers can reduce the number of employees needed (for example, Netflix Japan’s layoff of animators) (Jiang et al., 2023).
It isn’t just the output of AI models that can have an economic impact on artists, however. Harm, both financial and emotional, can occur based on whether people have – or even could have – consented to the content they created and shared in a certain context (Nissenbaum, 2004) (especially in the complex context of social media, e.g., an artist with a physical gallery presence whose work was photographed by a visitor) being repurposed by machine learning models and the people – and companies – who control or use them (Solon, 2019; Baio, 2022; Ho, 2024; the Midjourney team, 2023). The harm done by the AI model can be incipient before the AI model was ever conceived of; the complications of the model creation pipeline are discussed in more detail in Sec. 2.4.
Furthermore, the issue of copyright protection for artworks (made by humans) to train commercial AI models (without consent, attribution, or sufficient compensation to human authors) is a matter of concern for artists, creative industries, and federal agencies (Commission, 2023; Cho, 2023; Team, 2024; Patel and Palazzolo, 2024; David, 2024; Marcus and Southen, 2024).

### 14 May 2024 v4

@@ Abstract.
Generative AI tools are used to create art-like outputs and sometimes aid in the creative process. These tools have potential benefits for artists, but they also have the potential to harm the art workforce and infringe upon artistic and intellectual property rights. Without explicit consent from artists, Generative AI creators scrape artists’ digital work to train Generative AI models and produce art-like outputs at scale. These outputs are now being used to compete with human artists in the marketplace as well as being used by some artists in their generative processes to create art. We surveyed 459 artists to investigate the tension between artists’ opinions on Generative AI art’s potential utility and harm. This study surveys artists’ opinions on the utility and threat of Generative AI art models, fair practices in the disclosure of artistic works in AI art training models, ownership and rights of AI art derivatives, and fair compensation. Results show that a majority of artists believe creators should disclose what art is being used in AI training, that AI outputs should not belong to model creators, and express concerns about AI’s impact on the art workforce and who profits from their art. We hope the results of this work will further meaningful collaboration and alignment between the art community and Generative AI researchers and developers.
@@ 1. Introduction
Art has a rich history in co-evolution with the development of new technologies (Etro and Galenson, 2023). The advent of Artificial Intelligence (AI) tools used to create or assist in the creative process carries great potential utility for artists but also harbors the significant potential for harm to the art workforce and artistic and intellectual property. To examine this tension between utility and threat, we surveyed 459 artists about their views on AI-generated art. The survey results reveal the nuanced perspectives of artists on Generative AI. When the discussion is thorough and directly related to their art practice, it prevents their views from being reduced to a simplistic binary sentiment, which fails to represent the diversity of their field. By capturing artists’ views through a survey, this method moves beyond anecdotal evidence and provides a counterbalance to third-party opinions as well as those of the Generative AI industry.
@@
We summarize our research results as follows: (Q1) 61.87% of survey participants agreed that AI models are a threat to art workers. (Q2) 44.88% of survey participants agreed that AI art models are a positive development in the field of art (22.03% of survey participants agreed that both AI was a positive development and a threat). (Q3) Artists overwhelmingly agreed (80.17%) that model creators should be required to disclose in detail what art and images they use to train their AI models. (Q4) Artists were split between thinking that derivatives from AI art models should be owned by the original artists whose work is represented in the AI art (41.39%) and the user who used the model to create the AI art (39.22%). In our study, 26.80% of the participants agreed that the work and its derivatives should be considered the property of the AI model creators. We also looked into what types of compensation artists consider to be fair in exchange for the use of their artwork to train an AI art model. Half of the respondents (50.97%) stated that they did not need compensation, but what mattered to them was who was profiting from the use of their artwork. The most common fair compensation option selected by the survey participants was “I don’t need to profit, but I don’t want for-profit companies to profit from my art” (22.80%). This suggests that artists are not entirely against their art being used in AI art model training data but oppose for-profit companies benefiting financially.
@@ 2. Background
There is ongoing disagreement regarding whether the output of Generative AI can be considered art (US Federal Trade Commission, 2023; Jiang et al., 2023); we will not examine the topic in depth here. For convenience, we’ll refer to AI-generated output as “art” in this paper without asserting or refuting its artistic merit.
Jiang et al. define art as a human endeavor linked to culture and experience (Jiang et al., 2023). They argue that human creativity, distinguished by unique inspiration and corporeal engagement, differs from AI-generated work. Recent court decisions in the US have similarly denied copyright protection to autonomously produced AI creations, underscoring the role of human creativity in copyright law (US District Court for the District of Columbia, 2023).
Three compelling arguments against considering AI output as art emerge: art’s human origin devoid of ”organic pressure,” the intimate connection between artist and world, and art’s unique role in human experience (Jiang et al., 2023; Huyghe, 1970; Pfeifer and Bongard, 2006). Regardless of the artistry debate, AI-generated creations can meet commercial standards, like book covers, entering diverse artistic spheres, and impacting practitioners.
Artistic use of technology and AI tools (Cetinic and She, 2022) is not novel; artists historically adopt tools, including machines and automation, often innovating beyond standard applications (Huyghe, 1970). Artists engaging with technology question its role in art and explore its boundaries, raising queries about intentionality, distraction, and artistic agency (Karson, 2023).
@@
It is important to note that, although how technology should be incorporated into art may be up for debate, that technology is part of the artistic process is not new or ambiguous (Mazzone and Elgammal, 2019). Artists have always used tools (Huyghe, 1970), and many art forms involve machines (such as woodwork, metalwork, and digital art) and automation (such as digital printers and cameras). Some artists also choose to use technology in ways other than its intended purpose or develop new art-making techniques by breaking technology. Such artists may be both suspicious and intrigued by the technology they use, questioning the relationship between humans and machines and expanding art practices and processes. When presented with new technology, an artist may ask questions like “How can I keep my work with the machine intentional? How might this technology distract me? Is the machine using me, or am I using it? Is this image-making or art-making? How does this automation serve my artistic intentions and practice?” (Karson, 2023).
@@
In the case of machine learning, a crucial question for many artists is “What data was used to train this model?” Such inquiries are part of the artistic tradition of challenging the status quo and are in stark contrast to how the creators of Generative AI models have presented their tools for use to the general public.
2.1. Generative AI Models
Generative AI models can generate outputs across various media types based on user inputs, such as text prompts or multimedia (Gozalo-Brizuela and Garrido-Merchán, 2023). Key players in textual Generative AI include large language models (LLMs) like OpenAI’s ChatGPT and GPT-4, while visual models like StableDiffusion and Midjourney modify images using iterative noise adjustments guided by diffusion processes.
The field of Generative AI has sparked significant excitement due to its broad range of applications, including both anticipated ones valued in the tens of billions of dollars and substantial academic and industrial research efforts (Bubeck et al., 2023; Abdin et al., 2023; Jiang et al., 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2023). This expansive field encompasses thousands of models and tools that are driving research across various academic disciplines. However, alongside this enthusiasm, there is a pressing need for a more critical examination of the ethical implications associated with these tools and methods (Barnett, 2023). The deployment of these tools beyond academia has resulted in adverse consequences such as personal and financial losses, widespread information manipulation, and disruptions to sociotechnical infrastructure (Ferrara, 2024; Walker et al., 2023).
@@ 2.1. Generative AI Models
For the purposes of our paper, we define Generative AI as being a subset of tools or processes that combine massive probability distributions with sense-making algorithms in order to create output that, taken as a whole, was not included in the training data (Jiang et al., 2023). The probability distribution should be based on the training data. The sense-making algorithm could be based on several strategies, such as a physical process like diffusion or an architectural approach like transformers/neural nets, although in the end, it should produce something akin to a latent/embedding space; a place to hold meaning (Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023; Grand et al., 2022). Research is ongoing into the various distributions and processes that can be implemented in Generative AI models to achieve specific behaviors and results (Xu et al., 2023). However, irrespective of details, all of these models use advanced architectures in conjunction with large amounts of data gathered through brute-force methods across vast swathes of the publicly accessible internet (Xu et al., 2023; Zhao et al., 2023).
@@
In this paper, we are interested in how artists perceive Generative AI, what its impact on art-related careers might be, and what moral, ethical, artistic, and financial implications are coming hand-in-hand with the broader flourishing of Generative AI (Stanford Institute for Human-Centered Artificial Intelligence, 2023; Gozalo-Brizuela and Garrido-Merchán, 2023; Zhao et al., 2023). Therefore, we touch on three relevant topics – first, related academic work; second, the economic context for this discussion; and third, what sort of stages are within the pipeline of Generative AI models and how those stages can each cause harm (a specific and noteworthy type of impact) – within this Background section.
@@
Although in our paper we focus more on the pressing potential for harm, it is worth noting that there is also growing excitement around Generative AI’s capabilities and how it could be incorporated into artistic processes (Karson, 2023). For example, artist Cosmo Wenman worked with blind researchers Brandon Biggs, Lindsay Yazzolino, and Joshua Miele to create images using the Midjourney tool. Their team noted that programs like Midjourney can find “weird and thematic intersections” in art that go unnoticed by human eyes. They also discussed the role of chaos in creating “visual intrigue” and how technology can be used to “enhance agency.” Wenman called the new technology “essentially moveable type for visual rhetoric” (Schrader, 2023).
@@ 2.2. Related work
A paper by Jiang et al. (Jiang et al., 2023) analyzes Generative AI art models and their potential impact on the world through the lens of diverse academic sources (both philosophical and scientific). They describe many opportunities for harm as a result of such products: violation of the consent and privacy of artists (Ho, 2024), a lack of dataset transparency (shielding violations and biases) (Ho, 2024), a failure to uphold the contextual integrity of content (Nissenbaum, 2004), harm to the livelihood of the artist, biases in the training data, automating processes in a way that is detrimental to the creation of art, reduced accessibility, erasure of human works by obscurity, and increasing reluctance by artists to share their works openly.
The harms Jiang et al. (Jiang et al., 2023) foresee for artists fall into two categories, financial and emotional, which manifest as the replacement of a human artist on a project, the impersonation of an artist without consent (Gozalo-Brizuela and Garrido-Merchán, 2023), or plagiarism. Impersonation, in this context, could be weaponized for nefarious purposes (Ho, 2024). Moreover, these two kinds of harm, enacted against individuals at scale, create societal-level harm. Specific societal harms Jiang et al. discuss in their work include data laundering – through academic or open-source projects where public funds are funneled into research with private profits – the proliferation of stereotypes and “hegemonic views,” and erasure by obscurity (i.e., that AI-generated outputs can be created much more quickly than human ones). They also point out that US copyright and intellectual property jurisprudence needs to be better equipped to deal with these new tools and that open sourcing, often called on as a solution to software-related problems, is incomplete in this case (Jiang et al., 2023).
@@ 2.3. Economic context: livelihood of the artist
AI’s integration into the global economy is rapidly expanding, as evidenced by the increasing number of AI-related job postings across various economic sectors (Stanford Institute for Human-Centered Artificial Intelligence, 2023; Squicciarini and Nachtigall, 2021). Employers worldwide are seeking AI-skilled workers, with many companies opting to adopt AI tools each year. The primary benefit reported by these companies is a reduction in costs, often achieved through layoffs and the substitution of human roles with AI-driven solutions (Stanford Institute for Human-Centered Artificial Intelligence, 2023; Gmyrek et al., 2023; Alekseeva et al., 2021).
Some of these impacts are already coming to bear on the economic reality of artists (Chow, 2020; Henderson, 2010). Entry-level positions in film, TV, and gaming seem particularly at risk, and the conceptual period of a project seems to be a popular phase in which to replace artists with AI. Traditional niches like illustrating book covers can now be filled by the output of Generative AI, and major employers can reduce the number of employees needed (for example, Netflix Japan’s layoff of animators) (Jiang et al., 2023).
The economic impact on artists is not solely determined by the output of AI models. Harm, both financial and emotional(Caporusso et al., 2023), can arise from the use of content without explicit consent. This includes situations where individuals may not have anticipated their work being repurposed by machine learning models or the entities controlling and utilizing these technologies (Nissenbaum, 2004). For instance, consider an artist with a physical gallery presence whose work is photographed by a visitor and subsequently used in ways they did not foresee, especially in the intricate landscape of social media (Ho, 2024).
Moreover, the issue of copyright protection becomes paramount when AI models are trained using human artworks without proper consent, attribution, or adequate compensation for the original creators. This concern is shared among artists, creative industries, and government agencies alike (US Federal Trade Commission, 2023; Marcus and Southen, 2024).